Calculate the value of two resistance which when connect in series gives 50ohms and 8ohm when in parallel
1
Expert's answer
2020-11-10T06:57:42-0500

The resistance of serial resistors is:


"R = R_1 + R_2 = 50\Omega +8\Omega = 58\Omega"

The resistance of parallel resistors is:


"R = \dfrac{R_1R_2}{R_1+R_2} = \dfrac{50\Omega\times 8\Omega}{50\Omega+8\Omega}\approx 6.9\Omega"

Answer. In series: 58 ohms, parallel: 6.9 ohms.